# Optimal and Fair Encouragement Policy Evaluation and Learning

### v3

@@ 5 Additional Fairness Constraints and Policy Optimization
We assume that we are using doubly robust/orthogonalized estimation as in proposition 4 and, hence, state our results depending on the estimation error of nuisance vector $\eta$. The next theorem summarizes the out-of-sample performance of the two-stage algorithm of Algorithm 2, $\hat{Q}_{2}$.
@@
Suppose that the mean-squared error of the nuisance estimates is upper bounded w.p. $21-\delta/$ 2 by $\chi_{n,\delta}^{2}$ over the randomness of the nuisance sample: $\max_{l}\{\mathbb{E}[(\hat{\eta}_{l}-\eta_{l})^{2}]\}_{l\in[L]}\coloneqq\chi_{n}^{2}$
@@
The specific benefits of the two-stage approach are that 1) the constants are improved from being absolute, structure-agnostic bounds to depending on the variance of low-regret policies, which also reflects the improved variance from the use of doubly robust estimation as in proposition 4, and 2) it allows less conservative satisfaction of the fairness constraint out of sample.
@@ 6 Connections to IV Estimation
Nonadherence has classically been studied in econometrics with IVs. In this work, we focus primarily on the implications of Assumption 2. Common IV analyses impose additional assumptions such as monotonicity or no defiers; these can impose restrictions in behavior that empirical studies show are sometimes untrue. Nonetheless, in other domains, these assumptions can be more or less plausible, and so, in this section, we discuss connections to other IV models for completeness.
@@ 7.1.2 Model and descriptive statistics: Heterogeneous takeup and efficacy
Next, we include descriptive statistics to provide evidence on the existence of treatment effect and compliance heterogeneity and potential drivers of inequity in unconstrained optimal encouragement regimes by disaggregating the treatment effect and compliance heterogeneity (vs. the heterogeneity in the treatment effect of the nudge alone). In LABEL:tbl-regressionRX, we display the results of a logistic regression of recommendation on covariates. In the original data, significant explanatory factors include household size, whether the interview was conducted by phone or in person, gender (female), race (nonwhite), age, whether the beneficiary has children, whether the beneficiary receives the maximum benefit amount, and time lapse to the interview date. Next, in LABEL:tbl-regressionTRX, we include the regression table from a recommendation-interacted logistic regression of treatment $T$ on $(X,R,R\times X),$ i.e., covariates and interacted interview reminder. We find significant heterogeneity based on age (interview reminders are less efficacious as age increases), while reminder efficacy increases for those without income in the previous quarter (which indicates more need for benefits/a lower opportunity cost of interview attendance) and with years since first SNAP enrollment (which indicates greater familiarity with the recertification process).
In Figure 2, we compare the distributions within each race (binarized to white or nonwhite) of the heterogeneous treatment effect $\mathbb{E}[Y(1)-Y(0)\mid X],$ the heterogeneous compliance effect $p_{1\mid 1}(X)-p_{1\mid 0}(X),$ and the heterogeneous encouragement effect, which is the product of the previous two: $(\mathbb{E}[Y(1)-Y(0)\mid X])(p_{1\mid 1}(X)-p_{1\mid 0}(X))$. In a setting with fixed costs of recommendation and treatment, the default approach of viewing the recommendation as the treatment, i.e., an optimal unconstrained encouragement regime, sorts by the latter score.
In this setting, we find that treatment efficacy is broadly correlated with takeup: Higher treatment effects are rank correlated with higher compliance effects. However, the strength of this relationship varies across groups, leading to inequality. We quantify this via the Spearman’s $\rho$ rank correlation coefficient. It measures the correlation of the ranks of two random variables and hence indicates whether a monotonic (but not necessarily linear) relationship exists between them. Among white beneficiaries only, the rank correlation coefficient between heterogeneous treatment and compliance effects is 0.68, while it is 0.45 for nonwhite beneficiaries. While both correlations are statistically significantly positive, the relationship is weaker for nonwhite than for white beneficiaries, in particular with regard to those most likely to take up. This differential efficacy can lead to potential inequity.
A central concern here is that worse-off groups might have more to gain by accessing the treatment (i.e., have heterogeneous treatment effects of greater magnitude) but may respond less to nudges/recommendations into treatment (because of behavioral frictions, time/resource constraints, higher need, etc.). In this setting, decreasing marginal utilities of money establish the former, while the process of filling out the burdensome benefits application or making it to a recertification interview under time, childcare, transportation, or resource constraints speaks to the relative difficulty posed by the latter. The density plots in Figure 2 underscore this phenomenon. The first plot shows the treatment effect heterogeneity: The distribution is wider for nonwhite than for white beneficiaries; i.e., it has more mass at higher magnitudes of $\tau$ and less at lower magnitudes of $\tau$. However, we see the opposite effect on compliance: While the results are relatively weak, they are bimodal, with a substantially greater probability density at highercompliance treatment effects for white than for nonwhite beneficiaries. One reason for this could be the language barrier for non-English-speaking applicants.
@@ 7.1.3 Comparison of potential improvements upon the status quo
Next, we consider potential improvements from automatic enrollment in reminders (recommend to all) or, under potential budget constraints, budget-optimal targeted allocations. In Figure 1, we compare all of these potential policies. The y-axis measures different values of $E[Y(\pi)]$ in differently colored series of lines (red, purple, blue): red for the group $A=1$ (nonwhite), $E[Y(\pi)\mid A=1]$; purple for the entire population, $E[Y(\pi)]$; and blue for $A=0$ (white), $E[Y(\pi)\mid A=0]$. We optimize the policies on a 70% training split. The off-policy values are obtained via doubly robust estimation, with logistic regression as the propensity score and gradient-boosted regression on log-transformed outcomes for outcome regression. However, because of the zero-inflated/heavy-tailed outcomes, we find some instability in the off-policy estimates computed on the test set alone. Thus, we pool the data and obtain the plotted estimates of $E[Y(\pi)\mid A=1]$ from the entire dataset.
@@
Our final analysis considers equity in allocation efficiency, specifically the lift or improvements to decisions around who receives recommendations. In Figure 3, we evaluate the lift in different performance metrics conditional on protected group and conditional on the group’s being encouraged under a resource-constrained policy—the y-axes show $\mathbb{E}[c(Y\pi))\mid A=a,\pi(X)=1],\mathbb{E}[(Y(1)-Y(0))\mid A=a,\pi(X)=1],\mathbb{E}[(T(1)-T(0))\mid A=a,\pi(X)=1]$. The x-axis ranges over thresholds on the conditional encouragement effect, $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\times\tau(X,A)$. Thus, the x-axis explores the range of the population covered under the budget allocation, with the budget averaged over the full population on the left-hand side and encouragement restricted to few individuals on the right-hand side. This represents the range of resource-constrained rules under a naive ITT approach. We do find, for large ranges of moderate budgets, that the naive efficient allocation results in disparities in average outcomes for white and nonwhite beneficiaries. The next two figures on the right disaggregate the individual encouragement effect into the heterogeneous causal effect $\tau$ of interview attendance and the compliance effect $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))$ of text message reminders on interview attendance. Consistent with the earlier results on the strong positive rank correlation of $\tau$ and $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))$, both plots are nearly nondecreasing as we threshold based on increasing individual encouragement scores $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\times\tau$. However, as we also saw earlier, this correlation is weaker for nonwhite beneficiaries, and we can observe a small region of nonmonotonic behavior. Comparing these two plots, we see that the disparities in naive ITT-based allocation are in fact driven almost entirely by disparities in access, which would persist under a large range of budgets. Disaggregating our analysis by looking separately at treatment vs. compliance effects and at covariate-heterogeneous effects allows us to quantify access gaps that would persist even under efficient allocations.
@@
With these findings in mind, we turn to a final investigation: Suppose that we could consider allocations with different budgets per group rather than a global budget. Given that different compliance responses lead to unequal improvements, how much more outreach toward nonwhite populations would be needed under the allocation to equalize the amount of improvement over the no-allocation baseline? We analyze this question in Figure 4. We now examine policies that consider different thresholds for different groups, thereby resulting in different budgets, $\pi_{\%,A}=\mathbb{I}[(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)>q_{A}].$ Typically, $q_{A}$ is a quantile of the score distribution and therefore corresponds to some budget percentile. We let $F_{A}(p)=P(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)\leq p)$ denote the cumulative distribution function (CDF) of the heterogeneous encouragement score, and therefore, the groupwise budget at some threshold $q_{a}$ is $F^{-1}_{A}(q_{A})$. We now range over these thresholds separately for the different groups on the $x$-axis; the $y$-axis plots the average improvement in the allocated subgroup, $E[p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)\mid A,\pi_{\%,A}=1]$. In Table 2, we illustrate for a few potential groupwise budget percentages of $A=0$ what improvement is achieved for $A=0$ and what percentage of the budget would be required for $A=1$ to achieve equal improvement. The equivalent budget allocations required to target the same percentage of the group population can differ by 20 to 25 percentage points. At thresholds $q=F_{0}^{-1}(p)$ corresponding to percentiles $p=50\%,75\%$ and $90\%$ of the $A=0$ group, and therefore budgets of $1-p=50\%,25\%$ and $10\%$, we observe average improvements of \$$\$12,\$21.97$ and \$$\$39.37$ in the $A=0$ group, and it would require an equivalent budget of ${}72\%,50\%$ and $29.2\%$ of the $A=1$ population to achieve equal improvements, respectively. Given that 80% of individuals are nonwhite (i.e., the $A=1$ group is four times the size of the $A=0$ group), substantially different groupwise budgets (in terms of number of people covered by the budget allocation) would be required to equalize improvements.
@@
Figure 2 and Figure 3 illustrate that this inequity arises because of the differential effectiveness of the reminder in improving the probability of interview attendance, suggesting either that a design with further communication could be helpful or that other, structural compliance barriers could persist. Proceeding with the analysis of reminder efficacy, we find that while racial disparities in current patterns of self-selection into encouragement are small, self-selection is still inefficient for targeting: Those who would benefit the most from receiving reminders do not sign up for them. We assess budget-optimal targeted allocations that can improve upon the outcome under self-selection with as little as 10% of the budget. The heterogeneous effects of the reminder on interview completion, i.e., the differences that the nudge makes, are smaller for nonwhite beneficiaries, but this group generally receives higher benefits overall. Overall, any efficiency-targeted allocation rule improves groupwise outcomes (i.e., average 12-month benefits for white or nonwhite beneficiaries) relative to those in the no-reminder baseline.
@@
In Figure 6, we provide descriptive information illustrating heterogeneity (including by protected attribute) in adherence and effectiveness. We observe wide variation in judges’ assignment of supervised release beyond the recommendation. We use logistic regression to estimate outcome and treatment response models. The first figure shows our estimates of the causal effect by gender (with similar heterogeneity by race). The outcome is failure to appear, so negative scores are beneficial. The second figure illustrates the difference in responsiveness: how much more likely decision-makers are to assign treatment when there is than when there is not an algorithmic recommendation to do so. The last figure plots a logistic regression of the lift in responsiveness on the causal effect $\tau(x,a)=\mu_{1}(x,a)-\mu_{0}(x,a)$. We observe disparities in how responsive decision-makers are conditional on the same treatment effect efficacy. This is, importantly, not a claim of animus because decision-makers did not have access to causal effect estimates. Nonetheless, disparities persist.
@@
Figure 6: Policy value $V(\pi^{\lambda})$, treatment value $\mathbb{E}[T(\pi^{\lambda})\mid A=a],$ for $A$ = race, gender.
In Figure 6, we highlight results from constrained policy optimization. The first two plots in each set illustrate the objective function value and $A=a$ average treatment cost for $A=$ race (nonwhite/white) and gender (female/male), respectively. We use costs of $100$ for $Y=1$ (failure to appear, 00 for $Y=0$, and $20$ when $T=1$ (set arbitrarily)) and minimize costs. On the x-axis, we plot the penalty $\lambda$ that we use to assess the solutions of Proposition 9. The vertical dashed line indicates the solution achieving $\epsilon=0$, i.e., parity in treatment takeup. Near-optimal policies that reduce treatment disparity can be of interest given advocacy concerns about how the expansion of supervised release could increase the surveillance of already surveillance-burdened marginalized populations. We see that, indeed, for race, surveillance parity–constrained policies can substantially reduce disparities for nonwhite defendants while not increasing surveillance on white defendants that much: The red line decreases significantly with a low increase of the blue line (and low increases to the objective value). On the other hand, for gender, the opportunity for improvement in the surveillance disparity is much smaller.
@@ C.1 Additional discussion on constrained optimization
C.2 Additional discussion on Algorithm 2 (general algorithm)
@@ C.2.1 Weighted classification reduction and off-policy estimation
There is a well-known reduction of optimizing the zero-one loss for policy learning to weighted classification. Note that the reductions approach of [Agarwal et al., 2014] works with the Lagrangian relaxation which only further introduces datapoint-dependent additional weights. Notationally, in this section, for policy optimization, $\pi\in\{-1,+1\},T\in\{-1,+1\}$ (for notational convenience alone). We consider parameterized policy classes so that $\pi(x)=\pi(1\mid x)=\operatorname{sign}(f_{\beta}(x))$ for some index function $f$ depending on a parameter $\beta\in\mathbb{R}^{d}$. Consider the centered regret $J(\pi)=\mathbb{E}[Y(\pi)]-\frac{1}{2}\mathbb{E}[\mathbb{E}[Y\mid R=1,X]+\mathbb{E}[Y\mid R=0,X]].$ We summarize different estimation strategies via the score function $\psi_{(\cdot)}(O),$ where ()$(\cdot)\in\{DM,IPW,DR\}$: the necessary property is that ${\mathbb{E}[\psi\mid X]=\mathbb{E}[Y\mid R=1,X]-\mathbb{E}[Y\mid R=0,X]}$. The specific functional forms of these different estimators are as follows, where $\mu^{R}_{r}(X)=\mathbb{E}[Y\mid R=r,X]:$
@@ C.2.2 Additional fairness constraints and examples in this framework
In this section we discuss additional fairness constraints and how to formulate them in the generic framework. Much of this discussion is quite similar to [Agarwal et al., 2018] (including in notation) and is included in this appendix for completeness only. We only additionally provide novel identification results for another fairness measure on causal policies in Section C.2.3, concrete discussion of the reduction to weighted classification, and provide concrete descriptions of the causal fairness constraints in the more general framework.
@@
We let the conditioning events $\mathcal{E}_{a}=\{A=a\},\mathcal{E}_{\circ}=\{\text{True}\},$ i.e. conditioning on the latter is equivalent to evaluating the marginal expectation. Then we express Equation 4 as a set of equality constraints $h_{a}(\pi)=h_{\circ}(\pi),$ leading to pairs of inequality constraints,
@@
[…]

### v4

@@ 6 Additional Fairness Constraints and Policy Optimization
We assume that we are using doubly robust/orthogonalized estimation as in proposition˜5 and, hence, state our results depending on the estimation error of nuisance vector $\eta$. The next theorem summarizes the out-of-sample performance of the two-stage algorithm of Algorithm˜2, $\hat{Q}_{2}$.
@@
Suppose that the mean-squared error of the nuisance estimates is upper bounded w.p. $1-\delta/2$ by $\chi_{n,\delta}^{2}$ over the randomness of the nuisance sample: $\max_{l}\{\mathbb{E}[(\hat{\eta}_{l}-\eta_{l})^{2}]\}_{l\in[L]}\coloneqq\chi_{n}^{2}$
@@
The specific benefits of the two-stage approach are that 1) the constants are improved from being absolute, structure-agnostic bounds to depending on the variance of low-regret policies, which also reflects the improved variance from the use of doubly robust estimation as in proposition˜5, and 2) it allows less conservative satisfaction of the fairness constraint out of sample.
@@ 7 Connections to IV Estimation
Nonadherence has classically been studied in econometrics with IVs. In this work, we focus primarily on the implications of Assumption˜2. Common IV analyses impose additional assumptions such as monotonicity or no defiers; these can impose restrictions in behavior that empirical studies show are sometimes untrue. Nonetheless, in other domains, these assumptions can be more or less plausible, and so, in this section, we discuss connections to other IV models for completeness.
@@ 8.1.2 Model and descriptive statistics: Heterogeneous takeup and efficacy
Descriptive statistics provide evidence on the existence of treatment effect and compliance heterogeneity and potential drivers of inequity in unconstrained optimal encouragement regimes by disaggregating the treatment effect and compliance heterogeneity (vs. the heterogeneity in the treatment effect of the nudge alone). In Table˜3, we display the results of a logistic regression of recommendation on covariates. In the original data, significant explanatory factors include household size, whether the interview was conducted by phone or in person, gender (female), race (nonwhite), age, whether the beneficiary has children, whether the beneficiary receives the maximum benefit amount, and time lapse to the interview date. Next, in Table˜4, we include the regression table from a recommendation-interacted logistic regression of treatment $T$ on $(X,R,R\times X),$ i.e., covariates and interacted interview reminder. We find significant heterogeneity based on age (interview reminders are less efficacious as age increases), while reminder efficacy increases for those without income in the previous quarter (which indicates more need for benefits/a lower opportunity cost of interview attendance) and with years since first SNAP enrollment (which indicates greater familiarity with the recertification process).
@@ 8.1.3 Comparison of potential improvements upon the status quo
Next, we consider potential improvements from automatic enrollment in reminders (recommend to all) or, under potential budget constraints, budget-optimal targeted allocations. In Figure˜2, we compare all of these potential policies. The y-axis measures different values of $E[Y(\pi)]$ in differently colored series of lines (red, purple, blue): red for the group $A=1$ (nonwhite), $E[Y(\pi)\mid A=1]$; purple for the entire population, $E[Y(\pi)]$; and blue for $A=0$ (white), $E[Y(\pi)\mid A=0]$. We optimize the policies on a 70% training split. The off-policy values are obtained via doubly robust estimation, with logistic regression as the propensity score and gradient-boosted regression on log-transformed outcomes for outcome regression. However, because of the zero-inflated/heavy-tailed outcomes, we find some instability in the off-policy estimates computed on the test set alone. Thus, we pool the data and obtain the plotted estimates of $E[Y(\pi)\mid A=1]$ from the entire dataset.
@@
With these findings in mind, we turn to a final investigation: Suppose that we could consider allocations with different budgets per group rather than a global budget. Given that different compliance responses lead to unequal improvements, how much more outreach toward nonwhite populations would be needed under the allocation to equalize the amount of improvement over the no-allocation baseline? We analyze this question in Figure˜4. We now examine policies that consider different thresholds for different groups, thereby resulting in different budgets, $\pi_{\%,A}=\mathbb{I}[(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)>q_{A}].$ Typically, $q_{A}$ is a quantile of the score distribution and therefore corresponds to some budget percentile. We let $F_{A}(p)=P(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)\leq p)$ denote the cumulative distribution function (CDF) of the heterogeneous encouragement score, and therefore, the groupwise budget at some threshold $q_{a}$ is $F^{-1}_{A}(q_{A})$. We now range over these thresholds separately for the different groups on the $x$-axis; the $y$-axis plots the average improvement in the allocated subgroup, $E[p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)\mid A,\pi_{\%,A}=1]$. In Table˜2, we illustrate for a few potential groupwise budget percentages of $A=0$ what improvement is achieved for $A=0$ and what percentage of the budget would be required for $A=1$ to achieve equal improvement. The equivalent budget allocations required to target the same percentage of the group population can differ by 20 to 25 percentage points. At thresholds $q=F_{0}^{-1}(p)$ corresponding to percentiles $p=50\%,75\%$ and $90\%$ of the $A=0$ group, and therefore budgets of $1-p=50\%,25\%$ and $10\%$, we observe average improvements of $\mathdollar 12,\mathdollar 21.97$ and $\mathdollar 39.37$ in the $A=0$ group, and it would require an equivalent budget of $72\%,50\%$ and $29.2\%$ of the $A=1$ population to achieve equal improvements, respectively. Given that 80% of individuals are nonwhite (i.e., the $A=1$ group is four times the size of the $A=0$ group), substantially different groupwise budgets (in terms of number of people covered by the budget allocation) would be required to equalize improvements.
@@
Figure˜1 and Figure˜3 illustrate that this inequity arises because of the differential effectiveness of the reminder in improving the probability of interview attendance, suggesting either that a design with further communication could be helpful or that other, structural compliance barriers could persist. Proceeding with the analysis of reminder efficacy, we find that while racial disparities in current patterns of self-selection into encouragement are small, self-selection is still inefficient for targeting: Those who would benefit the most from receiving reminders do not sign up for them. We assess budget-optimal targeted allocations that can improve upon the outcome under self-selection with as little as 10% of the budget. The heterogeneous effects of the reminder on interview completion, i.e., the differences that the nudge makes, are smaller for nonwhite beneficiaries, but this group generally receives higher benefits overall. Overall, any efficiency-targeted allocation rule improves groupwise outcomes (i.e., average 12-month benefits for white or nonwhite beneficiaries) relative to those in the no-reminder baseline.
@@
In Figure˜6, we provide descriptive information illustrating heterogeneity (including by protected attribute) in adherence and effectiveness. We observe wide variation in judges’ assignment of supervised release beyond the recommendation. We use logistic regression to estimate outcome and treatment response models. The first figure shows our estimates of the causal effect by gender (with similar heterogeneity by race). The outcome is failure to appear, so negative scores are beneficial. The second figure illustrates the difference in responsiveness: how much more likely decision-makers are to assign treatment when there is than when there is not an algorithmic recommendation to do so. The last figure plots a logistic regression of the lift in responsiveness on the causal effect $\tau(x,a)=\mu_{1}(x,a)-\mu_{0}(x,a)$. We observe disparities in how responsive decision-makers are conditional on the same treatment effect efficacy. This is, importantly, not a claim of animus because decision-makers did not have access to causal effect estimates. Nonetheless, disparities persist.
@@
Figure 6: Policy value $V(\pi^{\lambda})$, treatment value $\mathbb{E}[T(\pi^{\lambda})\mid A=a],$ for AA = race, gender.
In Figure˜6, we highlight results from constrained policy optimization. The first two plots in each set illustrate the objective function value and $A=a$ average treatment cost for $A=$ race (nonwhite/white) and gender (female/male), respectively. We use costs of $100$ for $Y=1$ (failure to appear, 0 for $Y=0$, and $20$ when $T=1$ (set arbitrarily)) and minimize costs. On the x-axis, we plot the penalty $\lambda$ that we use to assess the solutions of Proposition˜10. The vertical dashed line indicates the solution achieving $\epsilon=0$, i.e., parity in treatment takeup. Near-optimal policies that reduce treatment disparity can be of interest given advocacy concerns about how the expansion of supervised release could increase the surveillance of already surveillance-burdened marginalized populations. We see that, indeed, for race, surveillance parity–constrained policies can substantially reduce disparities for nonwhite defendants while not increasing surveillance on white defendants that much: The red line decreases significantly with a low increase of the blue line (and low increases to the objective value). On the other hand, for gender, the opportunity for improvement in the surveillance disparity is much smaller.
@@ C.1 Additional discussion on constrained optimization
C.2 Additional discussion on Algorithm˜2 (general algorithm)
@@ C.2.1 Weighted classification reduction and off-policy estimation
There is a well-known reduction of optimizing the zero-one loss for policy learning to weighted classification. Note that the reductions approach of [Agarwal et al., 2014] works with the Lagrangian relaxation which only further introduces datapoint-dependent additional weights. Notationally, in this section, for policy optimization, $\pi\in\{-1,+1\},T\in\{-1,+1\}$ (for notational convenience alone). We consider parameterized policy classes so that $\pi(x)=\pi(1\mid x)=\operatorname{sign}(f_{\beta}(x))$ for some index function $f$ depending on a parameter $\beta\in\mathbb{R}^{d}$. Consider the centered regret $J(\pi)=\mathbb{E}[Y(\pi)]-\frac{1}{2}\mathbb{E}[\mathbb{E}[Y\mid R=1,X]+\mathbb{E}[Y\mid R=0,X]].$ We summarize different estimation strategies via the score function $\psi_{(\cdot)}(O),$ where $(\cdot)\in\{DM,IPW,DR\}$: the necessary property is that $\mathbb{E}[\psi\mid X]=\mathbb{E}[Y\mid R=1,X]-\mathbb{E}[Y\mid R=0,X]$}. The specific functional forms of these different estimators are as follows, where $\mu^{R}_{r}(X)=\mathbb{E}[Y\mid R=r,X]:$
@@ C.2.2 Additional fairness constraints and examples in this framework
In this section we discuss additional fairness constraints and how to formulate them in the generic framework. Much of this discussion is quite similar to [Agarwal et al., 2018] (including in notation) and is included in this appendix for completeness only. We only additionally provide novel identification results for another fairness measure on causal policies in Section˜C.2.3, concrete discussion of the reduction to weighted classification, and provide concrete descriptions of the causal fairness constraints in the more general framework.
@@
We let the conditioning events $\mathcal{E}_{a}=\{A=a\},\mathcal{E}_{\circ}=\{\text{True}\},$ i.e. conditioning on the latter is equivalent to evaluating the marginal expectation. Then we express Equation˜4 as a set of equality constraints $h_{a}(\pi)=h_{\circ}(\pi),$ leading to pairs of inequality constraints,
@@
The corresponding coefficients of $M$ over this enumeration over groups $(\mathcal{A})$ and epigraphical enforcement of equality $(\{+,-\})$ equation (1), gives $\mathcal{K}=\mathcal{A}\times\{+,-\}$ so that $M_{(a,+),a^{\prime}}=\mathbf{1}\left\{a^{\prime}=a\right\},M_{(a,+),\star}=-1$, $M_{(a,-),a^{\prime}}=-\mathbf{1}\left\{a^{\prime}=a\right\},M_{(a,-),\star}=1$, and $\mathbf{d}=\mathbf{0}$. Further we can relax equality to small amounts of constraint relaxation by instead setting $d_{k}>0$ for some (or all) $k$.